OpenStudy (anonymous):

What are the x-coordinates of the solutions to this system of equations? x2 + y2 = 100 y = x + 2

OpenStudy (campbell_st):

well substitute the 2nd equation into the 1st equation \[x^2 + (x + 2)^2 = 100\] which becomes \[x^2 + x^2 + 4x + 4 - 100 = 0\] simplify the equation then solve x by factoring... hope it helps

OpenStudy (anonymous):

i don't understand how to do that @campbell_st

OpenStudy (anonymous):

Do you not know how to factor or simplify the equation?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

so both?

OpenStudy (anonymous):

yes no to both

OpenStudy (campbell_st):

well the only other suggest I have, to find the point of intersection is to graph the 2 equations with some software and then find the points of intersection.

OpenStudy (campbell_st):

use this site, just enter the equations separately and you'll find the points of intersection. tps://www.desmos.com/calculator

OpenStudy (anonymous):

there are 2points of intersection though

OpenStudy (campbell_st):

thats correct, it will show both... just type y = x + 2 press enter then x^2 + y^2 = 100 it will graph both equations.. then zoom out so you see the points of intersection... then click on the points to get the coordinates hope it helps

OpenStudy (anonymous):

it does thank you!